### v2

@@ 2.3 Gate Fidelity Characterization
Figure 3: GST results. SPAM-free Pauli transfer matrices, process infidelities, diamond distances, stochastic Pauli errors and coherent Pauli errors for a resonant-KCQ $X(\pi/2)$ gate with 367-ns gate time, b resonant-KCQ $Z(\pi/2)$ gate with 70-ns gate time, c detuned-KCQ $X(\pi/2)$ gate with 418-ns gate time and d detuned-KCQ $Z(\pi/2)$ gate with 70-ns gate time. The uncertainty represents the standard deviation of the model fitting.
2.4 Noise Bias Characterization
@@
The quantum operations usually introduces additional noises, which may degrade the noise bias. Therefore, it is essential to benchmark the noise bias of bias-preserving gate sets on KCQ. Based on the stochastic Pauli error channel, we define the bit-flip error as $p_{\text{bit}}=p_{x}+p_{y}$, the phase-flip error as $p_{\text{ph}}=p_{z}$ and the noise bias as $\eta=p_{\text{ph}}/p_{\text{bit}}$. We present the first experimental application of DRB protocol on the KCQ to extract the noise structure of the $Z(\theta)$ gates with high precision[42, 46]. This protocol is also extensible to multi-qubit gate benchmarking by a recent protocol on CX dihedral groups[42].
@@
The single-qubit $\mathbb{D}_{8}$ dihedral group is generated by $X(\pi)$ and $Z(\pi/4)$ gates. The $X(\pi)$ gate for KCQ is trivially a $\pi$ phase shift of the reference, which can be applied virtually by adding an extra phase to the following control and readout pulses in a noise-free manner. This noise-free virtual $X(\pi)$ gate generates a bias-preserving $\mathbb{D}_{8}$ dihedral group with $Z(\pi/4)$ gates, allowing the extraction of bit-flip and phase-flip error separately. However, the noise-free property of the virtual $X(\pi)$ gate leads to underestimation of the extracted errors of $Z(\theta)$ rotations because the extracted errors are weighted averages of both noise-free virtual $X(\pi)$ gate and noisy $Z(\theta)$ rotations. To address such underestimation, scaling factors of 1.07 for $p_{\text{bit}}$ and 1.02 for $p_{\text{ph}}$ derived from numerical simulations are introduced (Suppl. Info. Section 6). As a result, by applying the DRB protocol, we can extract the bit-flip errors, phase-flip errors, and the noise bias of the bias-preserving gates precisely.
@@
Figure 4: DRB results. a The bit-flip $p_{\text{bit}}$ and phase-flip $p_{\text{ph}}$ error of resonant- and detuned-KCQ. The error bars are smaller than the marker size and therefore not plotted. b The noise bias and phase-flip error on the same plot with the fault-tolerant regime shaded in yellow. The error bars indicate the fitting uncertainty of the DRB survival probability.
We explore the noise structure of the resonant-KCQ and detuned-KCQ as a function of $Z(\theta)$ gate times, with results shown in Fig. 4a, where the uncertainty of the extracted errors is significantly smaller than that from GST. For both the resonant-KCQ and detuned-KCQ, the bit-flip error is much lower than the phase-flip error. The bit-flip error increases when the gate time is short because the strong gate pulses introduce leakage into the higher excited states, which increases the likelihood of tunneling between the two wells of the pseudo-potential. Conversely, the phase-flip error increases when the gate time is long due to the limited phase-flip lifetime. Interestingly, several local minima in the bit-flip error are observed, potentially linked to detailed leakage mechanisms and offering insights into optimal gate time determination.
For gate times longer than 60 ns, where the leakage mechanism is not dominant, the detuned-KCQ has a lower bit-flip error than the resonant-KCQ while maintaining the same phase-flip error, which directly demonstrates the advantage of the detuned-KCQ. Within a large range of gate times, we can achieve a noise bias as high as 100 for resonant-KCQ and 250 for detuned-KCQ with a reasonable phase-flip error rates below $2\%$, as shown in Fig. 4b. According to the previous work, with the tailored XZZX surface code, such noise bias can lead to 50% improvement of the error threshold with standard CX gates[19].
We proceed to assess the feasibility of using KCQ for QEC. Previous studies reported that the XZZX surface code has an error threshold of $p_{\text{ph}}<0.98\%\pm 0.05\%$ with a noise bias $\eta\geq 50$[19]. Accordingly, we show the phase-flip error for different noise bias $\eta$ in Fig. 4b, with the fault tolerant region shaded in yellow. Above all, the detuned-KCQ demonstrates performance well within the fault tolerant region when the $Z(\theta)$ gate duration is between 70 ns to 97 ns, as indicated by the four points inside the fault tolerant region.
3 Discussion
In summary, we have characterized and benchmarked high-fidelity quantum operations on a noise-biased KCQ within a scalable planar superconducting architecture. By introducing a novel detuned stabilization drive to enhance the bit-flip lifetime, we demonstrate the first implementation of quantum nondemolition readout and high-fidelity single-qubit universal quantum operations on a detuned-KCQ, achieving performance crossing the fault-tolerant threshold of the XZZX surface code[19]. Using GST, we measure SPAM-free gate fidelities of over 92% for the $X(\pi/2)$ gate and 99% for the bias-preserving $Z(\pi/2)$ gate, representing the highest performance reported in this field. The noise bias characteristic is implied in the enhanced bit-flip lifetime and directly examined through the DRB protocol. Remarkably, the detuned-KCQ exhibits a bit-flip lifetime of $T_{z}\approx 1.2$ ms, with a phase-flip lifetime comparable to the resonant-KCQ counterpart. This results in a noise bias as high as 250, which is crucial for surpassing the error threshold of the XZZX surface code[19]. The device’s performance is primarily limited by the relaxation rate of the SNAIL oscillator, which can be mitigated through material improvement[48]. Further performance gains can be achieved by speeding up gate operations via pulse shaping[49, 50] and through optimized circuit designs[37].

### v3

@@ Abstract
Figure 3: GST results. SPAM-free Pauli transfer matrices, process infidelities, diamond distances, stochastic Pauli errors and coherent Pauli errors for gates on resonant- and detuned-KCQs. a, resonant-KCQ $X(\pi/2)$ gate with 367-ns gate time. b, resonant-KCQ $Z(\pi/2)$ gate with 70-ns gate time. c, detuned-KCQ $X(\pi/2)$ gate with 418-ns gate time. d, detuned-KCQ $Z(\pi/2)$ gate with 70-ns gate time. The uncertainty represents the standard deviation of the model fitting, and each circuit in GST is repeated by 1024 shots.
The model violation of GST under completely positive, trace preserving (CPTP) (46) constraints is as small as $N_{\sigma}=14.51$, indicating the validity of the completely positive and trace preserving assumption. The SPAM-free estimations of the PTMs, errors, and fidelities for $X(\pi/2)$ and $Z(\pi/2)$ gates are shown in Fig. 3. For $Z(\pi/2)$ gates, the stochastic Pauli $X$ and $Y$ errors are significantly smaller than the stochastic Pauli $Z$ errors, confirming the bias preservation. Our KCQs exhibit low coherent errors, as evidenced by similar diamond distances and process infidelity. In fact, the effects of coherent errors are negligible after Pauli twirling (Suppl. Info. Section 7). According to the extracted process infidelities of the quantum gates, the process fidelities of $Z(\pi/2)$ and $X(\pi/2)$ gates for resonant-KCQ are 99.3% and 94.7%, and the detuned-KCQ has similar performance with the fidelities of 99.2% and 92.5%, respectively, signifying state-of-the-art performance (24, 23). The slightly lower gate fidelity of the $X(\pi/2)$ for detuned-KCQ is mainly due to the longer gate time.
Even though GST provides a comprehensive description of the noise structures of quantum operations on KCQs and confirms the noise-bias structure for bias-preserving gates, the uncertainty of extracted stochastic Pauli $X$ and $Y$ errors for bias-preserving gates is so large that a faithful estimation of them is not possible. This is because the Pauli $X$ and $Y$ errors are too small to be measured by GST with a shallow circuit depth of 128 at most. Therefore, we will apply the noise-biased dihedral randomized benchmarking method with significantly deeper circuits to accurately extract the rare Pauli $X$ and Pauli $Y$ errors.
Noise-Bias Dihedral Randomized Benchmarking
@@
Quantum operations usually introduce additional noise, which may degrade the noise bias. Therefore, the naive estimations of noise-bias from bit-flip and phase-flip time measurements (39, 23, 24) are not accurate, and it is essential to benchmark the noise-bias of bias-preserving operations of the KCQ. Based on the stochastic Pauli error channel, we define the bit-flip error as $p_{\text{bit}}=p_{x}+p_{y}$, the phase-flip error as $p_{\text{ph}}=p_{z}$, and the noise bias as $\eta=p_{\text{ph}}/p_{\text{bit}}$. The conventional randomized benchmarking method based on Clifford gates will mix the bit-flip and phase-flip errors, thereby preventing the estimation of noise bias  (47). Therefore, we present the first experimental application of the DRB protocol on the KCQ to extract the noise structure of the $Z(\theta)$ gates with high precision (41, 40) by applying a benchmarking circuit with a maximum depth over 2000. This protocol is also extensible to multi-qubit gate benchmarking by a recent protocol on CX dihedral groups (41).
@@
The noise-biased DRB protocol samples random operations from the single-qubit $\mathbb{D}_{8}$ dihedral group, which is generated by $X(\pi)$ and $Z(\pi/4)$ gates. The $X(\pi)$ gate for KCQ is trivially a $\pi$ phase shift of the reference, which can be applied virtually by adding an extra phase to the following control and readout pulses in a noise-free manner. This noise-free virtual $X(\pi)$ gate generates a bias-preserving $\mathbb{D}_{8}$ dihedral group with $Z(\pi/4)$ gates. Therefore, the DRB protocol avoids mixing the bit-flip and phase-flip errors, allowing them to be benchmarked separately. However, the noise-free property of the virtual $X(\pi)$ gate leads to an underestimation of the extracted errors of $Z(\theta)$ rotations because the extracted errors are weighted averages of both the noise-free virtual $X(\pi)$ gate and noisy $Z(\theta)$ rotations. Even though such underestimation is minor due to the low weight of the $X(\pi)$ gate in $\mathbb{D}_{8}$, we proceed to address this underestimation by introducing a scaling factors of 1.07 for $p_{\text{bit}}$ and 1.02 for $p_{\text{ph}}$ calculated from numerical simulations (Suppl. Info., Section 6). As a result, by applying the DRB protocol, we can precisely extract the bit-flip errors, phase-flip errors, and noise bias of the bias-preserving gates.
@@
Figure 4: DRB results. a, The bit-flip $p_{\text{bit}}$ and phase-flip $p_{\text{ph}}$ errors of resonant- and detuned-KCQ, and corresponding estimation $\tilde{p}_{\text{ph}}$ and $\tilde{p}_{\text{bit}}$ based on the bit-flip and phase-flip times measured in Fig. 2. The error bars are smaller than the marker size and therefore not plotted. b, The noise bias as a function of phase-flip errors.
We presently explore the noise structure of the resonant-KCQ and detuned-KCQ as a function of $Z(\theta)$ gate times, with results shown in Fig. 4a, where the uncertainty of the extracted errors is significantly smaller than that from GST. We also plot the bit-flip and phase-flip errors as predicted by the bit-flip and phase-flip times based on the stochastic Pauli error model  (48). For both the resonant-KCQ and detuned-KCQ, the bit-flip error is much lower than the phase-flip error. Clearly, while the phase-flip errors align with the rates inferred from the phase-flip time (green line), the bit-flip error is typically much higher than predicted by the bit-flip time (red and blue lines for resonant- and detuned-KCQs, respectively). We attribute the high bit-flip error at short gate times to the leakage into higher excited states caused by the strong gate drive, which tunnels between the wells of the pseudo-potential. As the gate time increases, the bit-flip error approaches the error set by the bit-flip time. These observations highlight an underestimation of the bit-flip errors, and accordingly, an overestimation of the noise bias, based solely on the bit-flip and phase-flip times, underscoring the need for direct benchmarking of the operational noise characteristics.
As shown in Fig. 4a, for gate times longer than 60 ns, the detuned-KCQ has a lower bit-flip error than the resonant-KCQ while maintaining the same phase-flip error, indicating the advantage of the detuned-KCQ with higher noise bias. Fig. 4b further shows the noise-bias of both resonant- and detuned-KCQs with respect to their phase-flip errors. Using the detuned-KCQ, we can achieve a noise bias as high as 250 with phase-flip error rates below $2\%$ for the bias-preserving gate set. This is over twice as large as that for the resonant-KCQ without increasing the phase-flip error. Such high-fidelity bias-preserving gates with a large noise bias, together with high-fidelity quantum nondemolition readout and state initialization, represent the state-of-the-art performance and position detuned-KCQ as a promising platform for the future implementation of hardware-efficient quantum error correction.
Discussion
In summary, we have developed a complete quantum control toolbox for detuned-KCQs, including initialization, universal single-qubit gates, and quantum nondemolition readout. Moreover, we provide the first comprehensive experimental benchmarking of noise-biased qubit operations, confirming their bias-preserving properties and revealing the inaccuracy of noise bias estimation from idling bit-flip and phase-flip times. By introducing a detuned stabilization drive to enhance the bit-flip time, we demonstrate the first implementation of high-fidelity single-qubit universal quantum operations on a detuned-KCQ, achieving a bit-flip time of 1.2 ms, which is three times the conventional resonant-KCQ, without further introducing more phase-flip noises. Using GST, we extract the detailed noise structure of the noise-biased qubit operations, and measured SPAM-free gate fidelities of over 92.5% for the $X(\pi/2)$ gate and 99.2% for the bias-preserving $Z(\pi/2)$ gate, representing the highest performance reported in this field. The minimal bit-flip error and the high noise bias of the bias-preserving gates are further accurately measured by the DRB protocol. While the resonant-KCQ already shows a large noise bias as high as 100 with a phase-flip error rate below 2%, the detuned-KCQ has an even larger noise bias of 250 with a similar phase-flip error rate. With the DRB protocol, we confirmed the enhanced noise-bias without degraded phase-flip error for bias-preserving gates of detuned-KCQ, demonstrating the advantage of detuned-KCQ.
The decay rate of the SNAIL nonlinear oscillator primarily constrains the performance of the detuned-KCQ. Continued advancements in superconducting circuit materials are expected to reduce this decay rate (49), and faster gate operations enabled by pulse shaping (50, 51) and optimized circuit designs (33) may further enhance detuned-KCQ performance. In addition, the minor model violation observed in GST under the CPTP framework for noise characterization implies low non-trace-preserving errors, such as leakage, but a more direct and precise characterization of these errors would be beneficial for optimizing performance.